If point M is a midpoint of segment XY, and segment MX=(2x+3) and segment XY=(4x+6), find x...

Please Help............. It's either I think fraction 5/4 or number 4....
1
Expert's answer
2012-11-22T08:05:54-0500
X______M______Y
MX=XY/2 because M is a midpoint. So we can construct the following equation by substituting given expressions:
2x+3=(4x+6)/2
2x+3=2x+3
0=0 this equation holds for any value of x.
Thus, we can't find x from data available. The statement is insufficient.